[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu May 25, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu May 25 17:45:51 UTC 2017


HSU Shift Log for Thu May 25 2017 13:45:45 GMT-0400 (EDT)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues: GOES-16 imagery is still missing tiles in D2D. We've  
been told this will not be fixed until our method of ingest for AWIPS2  
GOES-16 imagery is changed from scp to LDM.

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170525/cfe54c3a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list